### v1

# Toward a generalization of Lehmer’s problem to adelic curves

Abstract.
In this short note, we investigate the generalization of Lehmer’s problem to finitely generated fields over $\mathbb{Q}$.
Keywords: Lehmer’s problem; adelic curves; heights
MSC: 11G50, 14G40.
1. Introduction
Moriwaki [12] has developed an Arakelov height theory for varieties over a finitely generated extension of a number field. He applied his theory to the study of Bogomolov problem over finitely generated fields over $\mathbb{Q}$ and obtained generalizations of results due to Ullmo [13] and Zhang [14]. As a corollary, he covered the original Raynaud’s theorem (Manin-Mumford’s conjecture). This Arakelov height theory is viewed as an example of an arithmetic intersection theory over an adelic curve in the sense of Chen and Moriwaki [7]. Burgos, Philippon and Sombra [6] have showed that the height of a variety over a finitely generated extension of $\mathbb{Q}$ can be written as an integral of local heights over the set of places of the field. Using their previous work on toric varieties in [5], they gave a combinatorial formula for the height of toric varieties over finitely generated fields over $\mathbb{Q}$.
This work is an attempt to generalize Lehmer’s problem to adelic curves. We show that Lehmer’s problem generalizes to a natural adelic curve defined on the field of rational functions $\mathbb{Q}(T_{1},\ldots,T_{d})$.
Let $f$ be a nonzero polynomial with complex coefficients such that $f(x)=a\prod_{i}(x-\alpha_{i})$. The Mahler measure of $f$ is defined as
(1) $m(f)=\log|a|+\sum_{i}\log\max(1,|\alpha_{i}|).$
If $f$ has integer coefficients then (1) shows that $m(f)$ is nonnegative. A classical theorem of Kronecker asserts that $m(f)=0$ if and only if $|a|=1$ and $\alpha_{i}$ are zero or roots of the unity.
Lehmer asked whether there exists a constant $c>1$ such that for every polynomial $f\in\mathbb{Z}[x]$ with $m(f)>0$, then $m(f)\geq c$. Lehmer’s question has not yet been settled. The best result in the direction of solving Lehmer’s problem is Dobrowolski’s result [10] which gives a lower bound depending on the degree of $f$.
Let $h:\overline{\mathbb{Q}}\rightarrow\mathbb{R}$ be the absolute logarithmic height. Then, Lehmer’s problem can be reformulated as follows. Let $\alpha$ be an algebraic number of degree $d$ over $\mathbb{Q}$ and suppose that it is not a root of unity, then there exists an absolute constant $c>0$ such that
$h(\alpha)\geq\frac{c}{d}.$
Analogues of this conjecture have considered in other contexts, see [9, 8, 11].
Let us point out that generalizations of Lehmer’s problem to higher dimensional varieties have been investigated by Amoroso and David, see for instance [1, 2, 3].
Chen and Moriwaki [7] have introduced the theory of adelic curves. For example, a number field can be viewed as an example of adelic curves. More interestingly, they developed an arithmetic intersection theory on these curves. For instance, one can define a height function on the set of points of adelic curves.
In this paper, we show that Lehmer’s problem generalizes to a natural adelic curve defined on the field of rational functions with coefficients in $\mathbb{Q}$. Let us denote by $S$ the adelic curve on $\mathbb{Q}(T_{1},\ldots,T_{d})$ defined in Section 2.2. In Theorem 4.1, we establish a generalization of Kronecker’s theorem to $S$.
$\{\alpha\in(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
where $h_{S}(\alpha)$ denotes is the height of the point. $[1:\alpha]$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})$, see Paragraph 2.1 or [12, Paragraph 3.2]. $h_{S}$ is called naive height function in [12].
Assuming Lehmer’s conjecture, then for every $\alpha$ in an algebraic closure of $\mathbb{Q}(T_{1},\ldots,T_{d})$ which is not a root of unity,
$h_{S}(\alpha)\geq\frac{c}{\deg(\alpha)},$
where $h_{S}$ is the standard height function (see (1)), $\deg(\alpha)=[\mathbb{Q}(T_{1},\ldots,T_{d})(\alpha):\mathbb{Q}(T_{1},\ldots,T_{d})]$ and $c$ is an absolute constant, see Theorem 4.3.
2. Adelic curves
In this section, we briefly recall the theory of adelic curves. Let $K$ be a commutative ring and $M_{K}$ be the set of all absolute values on $K$. Following [7], an adelic structure on $K$ consists of a measure space $(\Omega,\mathcal{A},\nu)$ equipped with a map $\phi:\omega\mapsto|\cdot|_{\omega}$ from $\Omega$ to $M_{K}$ satisfying the following properties:
  1. (i)
    $\mathcal{A}$ is a $\sigma$-algebra on $\Omega$ and $\nu$ is a measure on $(\Omega,\mathcal{A})$;
  2. (ii)
    for any $a\in K^{\ast}:=K\setminus\{0\}$, the function $\omega\mapsto\log|a|_{\omega}$ is $\mathcal{A}$-measurable, integrable with respect to the measure $\nu$.
This data is denoted by $S=(K,(\Omega,\mathcal{A},\nu),\varphi)$.
$S$ is said to be proper if the equality
$\int_{\Omega}\log|a|_{\omega}d\nu(\omega)=0,$
holds for each $a\in K^{\ast}$.
From now on, $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is a proper adelic curve and $K^{ac}$ an algebraic closure of $K$. Let $S\otimes_{K}K^{ac}=({K^{\mathrm{ac}}},(\Omega_{K^{\mathrm{ac}}},\mathcal{A}_{K^{\mathrm{ac}}},\nu_{K^{\mathrm{ac}}}),\phi_{K^{\mathrm{ac}}})$ be the algebraic extension of $S$ by ${K^{\mathrm{ac}}}$ (see [7, Paragraph 3.4.2] for the construction).
2.1. Height of points in $\mathbb{P}^{n}(K^{\mathrm{ac}})$
For $[a_{0}:\ldots:a_{n}]\in\mathbb{P}^{n}({K^{\mathrm{ac}}})$, the height of $[a_{0}:\ldots:a_{n}]$ with respect to the adelic curve $S$ is given as follows.
$h_{S}([a_{0}:\ldots:a_{n}])=\int_{\Omega^{ac}}\log\max(|a_{0}|_{\chi},\ldots,|a_{n}|_{\chi})\nu_{K^{\mathrm{ac}}}(d\chi),$
see [7, Paragraph 3.5] for more details about the construction.
For every $\alpha\in K^{\textrm{ac}}$, we set
(1) $h_{S}(\alpha):=h_{S}([1:\alpha]).$
Lemma 2.1.
The following set
$\{\alpha\in(K^{\mathrm{ac}})^{\ast}\mid h_{S}(\alpha)=0\},$
is a subgroup of $(K^{\mathrm{ac}})^{\ast}$.
Proof.
This follows from the following inequality
$\max(1,xy)\leq\max(1,x)\max(1,y),\quad\forall x,y\geq 0.$
and the product formula (i.e. the properness of $S$).
2.2. The adelic structure on $\mathbb{Q}(T_{1},\ldots,T_{d})$
Let $K=\mathbb{Q}(T_{1},\ldots,T_{d})$ be the field of rational functions of $d$ variables $T_{1},\ldots,T_{d}$ and with coefficients in $\mathbb{Q}$. $\mathbb{Q}(T_{1},\ldots,T_{d})$ can be endowed with an interesting adelic structure, see [7, Paragraph 3.2.6], or [12, Section 3] and [6, Example 2.5] for a more explicit construction. Let us review in details the case when $d=1$.
Let $\Omega=\Omega_{h}\coprod\mathcal{P}\coprod[0,1]_{\ast}$, where $\Omega_{h}$ is the set of all closed points of $\mathbb{P}_{\mathbb{Q}}^{1}\setminus\{\infty\}$, and $\mathcal{P}$ denotes the set all prime numbers, and $[0,1]_{\ast}$ the set of $t\in[0,1]$ such that $e^{2\pi it}$ is transcendental.
Let $\phi:\Omega\rightarrow M_{K}$ be the map sending $\omega\in\Omega$ to $|\cdot|_{\omega}$, defined as follows. Let $x\in\Omega_{h}$. $x$ defines a discrete valuation on $K$ which is denoted by $\mathrm{ord}_{x}(\cdot)$. Let $F_{x}$ be an irreducible polynomial in $\mathbb{Z}[T]$ with $F_{x}(x)=0$. Let $H(x)$ be the Mahler measure of the polynomial $F_{x}$. It is defined as follows.
$H(x):=\exp\left(\int_{0}^{1}\log|F_{x}(e^{2\pi it})|dt\right).$
Let $|\cdot|_{x}$ be the absolute value on $\mathbb{Q}(T)$ such that
$|\varphi|_{x}=H(x)^{-\mathrm{ord}_{x}(\varphi)},\quad\forall\varphi\in\mathbb{Q}(T).$
For $p$ a prime number, let $|\cdot|_{p}$ be $p$-adic value on $\mathbb{Q}$. It extends to $\mathbb{Q}[T]$ as follows.
$|\varphi|_{p}:=\max_{j=0,\ldots,d}|a_{j}|_{p},\quad\forall\varphi=\sum_{i=0}^{d}a_{i}T^{i}\in\mathbb{Q}[T],$
and then, by multiplicativity, to $\mathbb{Q}(T)$.
For any $t\in[0,1]_{\ast}$, let $|\cdot|_{t}$ be the absolute value on $\mathbb{Q}(T)$ such that
$|\varphi|_{t}:=|\varphi(e^{2\pi it})|,\quad\forall\varphi\in\mathbb{Q}(T),$
where ||$\cdot$|| ⋅ | denotes the usual absolute value of $\mathbb{C}$.
$\Omega_{h}$ and $\mathcal{P}$ are equipped with the discrete $\sigma$-algebras, and $[0,1]_{\ast}$ with the restriction of the Borel $\sigma$-algebra on $[0,1]$. Let $\mathcal{A}$ be the $\sigma$-algebra on $\Omega$ generated by the above $\sigma$-algebras on $\Omega_{h},\mathcal{P}$ and $[0,1]_{\ast}$ respectively. Let $\nu$ be the measure on $\Omega$ such that $\nu(\{x\})=1$ for $x\in\Omega_{h}$, that $\nu(\{p\})=1$ for any prime number $p$ and that the restriction of $\nu$ on $[0,1]_{\ast}$ coincides with the Lebesgue measure.
This data defines an adelic structure on $\mathbb{Q}(T)$. Moreover, $(\mathbb{Q}(T),(\Omega,\mathcal{A},\nu),\phi)$ is proper, see [7, Paragraph 3.2.5].
3. On Lehmer’s problem on adelic curves
Question 3.1.
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve and $K^{ac}$ an algebraic closure of $K$. Is there a constant $c>0$ such for every $\alpha\in K^{\mathrm{ac}}$ with $h_{S}(\alpha)>0$, then
$h_{S}(\alpha)\geq\frac{c}{[K(\alpha):K]}.$
When $K$ is $\mathbb{Q}$ and $S$ is the classical adelic curve associated with $\mathbb{Q}$, then this is the classical Lehmer’s problem.
Next, we show that the classical Lehmer’s problem generalizes naturally to the adelic curve $\mathbb{Q}(T)$.
4. Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$
Let $S$ be the adelic curve on $\mathbb{Q}(T)$ recalled in Paragraph 2.2. Let
$[\varphi:\psi]\in\mathbb{P}^{1}\left(\overline{\mathbb{Q}(T)}\right).$
We may assume that $\varphi$ and $\psi$ are integral over $\mathbb{Z}[T]$. Then, there exist $d,e\in\mathbb{N}$ and $f_{1}(T),\ldots f_{d}(T)$ and $g_{1}(T),\ldots,g_{e}(T)$ in $\mathbb{Z}[T]$ such that
$\varphi^{d}+f_{1}(T)\varphi^{d-1}+\ldots+f_{d-1}(T)\varphi+f_{d}(T)=0,$
and
$\psi^{e}+g_{1}(T)\varphi^{e-1}+\ldots+g_{e-1}(T)\psi+g_{e}(T)=0.$
Let
$c_{\varphi,\psi}:=\gcd(c(f_{d})^{e},c(g_{e})^{d}).$
where $c(f_{d})$ (resp. $c(g_{e})$) is the common divisor of the coefficients of $f_{d}$ (resp. $g_{e}$).
If $\{x\in\overline{\mathbb{Q}}\mid g_{e}(x)=f_{d}(x)=0\}$ is not empty, we let
$F(T):=\prod_{\begin{subarray}{c}x\in\overline{\mathbb{Q}}\\ g_{e}(x)=f_{d}(x)=0\end{subarray}}F_{x}(T)^{\min(\mathrm{ord}_{x}(f_{d}^{e}),\mathrm{ord}_{x}(g_{e}^{d}))}\,,$
with $F_{x}\in\mathbb{Z}[T]$ with coprime coefficients, irreducible in $\mathbb{Q}[T]$ and $F_{x}(x)=0$ for $x\in\overline{\mathbb{Q}}$. Otherwise, we let
$F(T)=1.$
By the product formula, we have
$h_{S}([\varphi:\psi])=h_{S}\Bigl{(}\Bigl{[}\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}:\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{]}\Bigr{)}.$
Let $K^{\prime}$ be a finite extension of $\mathbb{Q}(T)$ such that $\varphi$ and $\psi$ are in $K^{\prime}$.
Let $\omega\in\Omega_{K^{\prime}}$ with $\omega\mid x$ where $x\in\Omega_{h}$. If $x$ is a common zero of $f_{e}$ and $g_{d}$, then
$\begin{split}\log\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}=&\frac{1}{d}\log|f_{d}|_{x}-\frac{1}{de}\log|F|_{x}\\ =&-\frac{\mathrm{ord}_{x}(f_{d}(T))}{d}\log H(x)+\frac{\min(\mathrm{ord}_{x}(f_{d}(T)^{e}),\mathrm{ord}_{x}(g_{e}(T)^{d}))}{de}\log H(x).\end{split}$
and
$\begin{split}\log\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{d}}}\right|_{\omega}=&\frac{1}{e}\log|g_{e}|_{x}-\frac{1}{de}\log|F|_{x}\\ =&-\frac{\mathrm{ord}_{x}(g_{e}(T))}{e}\log H(x)+\frac{\min(\mathrm{ord}_{x}(f_{d}(T)^{e}),\mathrm{ord}_{x}(g_{e}(T)^{d}))}{de}\log H(x).\end{split}$
We conclude that
$\max\left(\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega},\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}\right)=1$
This equation holds if $x$ is not a common zero of $f_{e}$ and $g_{d}$.
Let $\omega\in\Omega$ with $\omega\mid p$ and $p\in\mathcal{P}$. We have
$\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}=\frac{|f_{d}(T)|_{p}^{\frac{1}{d}}}{|(c_{\varphi,\psi})^{\frac{1}{de}}|_{p}}=\left(\frac{|c(f_{d})^{e}|_{p}}{|c_{\varphi,\psi}|_{p}}\right)^{\frac{1}{de}},$
and
$\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}=\frac{|g_{e}(T)|_{p}^{\frac{1}{e}}}{|(c_{\varphi,\psi})^{\frac{1}{de}}|_{p}}=\left(\frac{|c(g_{e})^{d}|_{p}}{|c_{\varphi,\psi}|_{p}}\right)^{\frac{1}{de}}.$
We conclude that
$\max\left(\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega},\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}\right)=1.$
Case of $w\in\Omega$ with $w\mid t$ and $t\in[0,1]_{\ast}$. Note that $K^{\prime}_{\omega}=\mathbb{Q}(T)_{t}=\mathbb{C}$. We have then
$\begin{split}\int_{\begin{subarray}{c}w\in\Omega\\ w\mid t,\ t\in[0,1]_{\ast}\end{subarray}}\log\max\Bigl{(}&\Bigl{|}\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega},\Bigl{|}\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}\Bigr{)}d\nu(\omega)\\ =&\int_{t\in[0,1]_{\ast}}\frac{1}{[K^{\prime}:\mathbb{Q}(T)]}\left(\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\max\Bigl{(}\Bigl{|}\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega},\Bigl{|}\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}\Bigr{)}\right)dt.\end{split}$
Gathering the above computations, we obtain that
$\frac{1}{[K^{\prime}:\mathbb{Q}(T)]}\max\Bigl{(}\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl{|}\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}dt,\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl{|}\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}dt\Bigr{)}\leq h_{S}([\varphi:\psi]).$
Using [7, Lemma 3.5.4], we obtain
$\begin{split}\tfrac{1}{[K^{\prime}:\mathbb{Q}(T)]}\max\Bigl{(}\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{\omega}dt,\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}&\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{\omega}dt\Bigr{)}\\ &\leq[K^{\prime}:\mathbb{Q}(T)]h_{S}([\varphi:\psi]).\end{split}$
Hence
(2) $\begin{split}\max\Bigl{(}\int_{t\in[0,1]}\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{t}dt,\int_{t\in[0,1]}&\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{t}dt\Bigr{)}\\ &\leq[K^{\prime}:\mathbb{Q}(T)]h_{S}([\varphi:\psi]).\end{split}$
Theorem 4.1.
(A Kronecker’s theorem)
$\{\alpha\in(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
Proof.
For simplicity, we prove the theorem when $d=1$. Let $\varphi$ and $\psi$ be two coprime polynomials in $\mathbb{Z}[T]$. We have
$h_{S}([\varphi:\psi])=\int_{0}^{1}\log\max(|\varphi(e^{2\pi it})|,|\psi(e^{2\pi it})|)dt.$
Let us assume that $h_{S}([\varphi:\psi])=0$. This implies that
$\varphi(z),\psi(z),\ \text{and}\ \varphi(z)+y\psi(z)$
are generalized cyclotomic polynomials in the variables $y,z$ (see for instance [4]). We have used the following fact:
$\int_{0}^{1}\log\max(|\varphi(e^{2\pi it})|,|\psi(e^{2\pi it})|)dt=\int_{0}^{1}\int_{0}^{1}\log|\varphi(e^{2\pi it})+e^{2\pi is}\psi(e^{2\pi it})|dtds.$
It is not difficult to show that $\phi$ and $\psi$ are, in fact, roots of unity in $\overline{\mathbb{Q}}$.
Let $\alpha\in\overline{\mathbb{Q}(T)}$ with $h_{S}(\alpha)=0$. Using (2), we can conclude that $\alpha$ is a root of unity in $\overline{\mathbb{Q}}$.
Let $\varphi\in\overline{\mathbb{Q}(T)}$. Let $d:=[\mathbb{Q}(T)(\varphi):\mathbb{Q}(T)]$. We can assume
$[1:\varphi]=[\varphi_{1}:\varphi_{2}],$
with $\varphi_{1}\in\mathbb{Z}[T]$ and $\varphi_{2}\in\overline{\mathbb{Z}[T]}$ (integral closure of $\mathbb{Z}[T]$).
From (2) and assuming that Lehmer’s problem is true, we get
$h_{S}(\varphi)(=h_{S}([\varphi_{1}:\varphi_{2}]))\geq\frac{c}{\deg(\varphi)},$
where $\deg\varphi=[\mathbb{Q}(T)(\alpha):\mathbb{Q}(T)]$ and $c$ is the constant in Lehmer’s problem.
Remark 4.2.
If $\alpha\in\overline{\mathbb{Q}}$, then
$h_{S}(\alpha)=h(\alpha),$
where $h$ is the standard Weil’s height of $\alpha$.
We conclude that
Theorem 4.3.
Lehmer’s problem is equivalent to Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$.

### v2

@@
 He applied his theory to the study of Bogomolov problem over finitely generated fields over $\mathbb{Q}$ and obtained generalizations of results due to Ullmo [27] and Zhang [29].
-As a corollary, he covered the original Raynaud’s theorem (Manin-Mumford’s conjecture).
-This Arakelov height theory is viewed as an example of an arithmetic intersection theory over an adelic curve in the sense of Chen and Moriwaki [7].
+As a corollary, he recovered the original Raynaud’s theorem (Manin-Mumford’s conjecture).
+Using the height machinery developed in [24], Vojta [28] extended the classic Roth’s theorem to finitely generated fields.
 Burgos, Philippon and Sombra [9] have showed that the height of a variety over a finitely generated extension of $\mathbb{Q}$ can be written as an integral of local heights over the set of places of the field.
 Using their previous work on toric varieties in [8], they gave a combinatorial formula for the height of toric varieties over finitely generated fields over $\mathbb{Q}$.
-This work is an attempt to generalize Lehmer’s problem to adelic curves.
-We show that Lehmer’s problem generalizes to a natural adelic curve defined on the field of rational functions $\mathbb{Q}(T_{1},\ldots,T_{d})$.
-Let $f$ be a nonzero polynomial with complex coefficients such that $f(x)=a\prod_{i}(x-\alpha_{i})$. The Mahler measure of $f$ is defined as
+Chen and Moriwaki [12] have introduced the theory of adelic curves.
+For example, a number field can be viewed as an example of adelic curve.
+They developed an arithmetic intersection theory on these curves.
+For instance, one can define a height function on the set of points of adelic curves.
+The Arakelov height theory of [24] is then seen as an arithmetic intersection theory in the sense of [12].
+We briefly recall the classical Lehmer’s problem (for details, see §2).
+It can be formulated as follows: There exists a constant $c>0$ such that for every non zero algebraic number $\alpha$ of degree $d$ over $\mathbb{Q}$ that is not a root of unity, we have
+$h(\alpha)\geq\frac{c}{d}.$
+Lehmer’s question has not yet been settled. The best result in the direction of solving Lehmer’s problem is Dobrowolski’s result [18] (see (2)).
+Our work is an attempt to generalize Lehmer’s problem to adelic curves.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve and $K^{ac}$ an algebraic closure of $K$ (see § 4).
+A natural question arises: Is there a constant $c>0$ such for every $\alpha\in K^{\mathrm{ac}}$ with $h_{S}(\alpha)>0$, we have
+$h_{S}(\alpha)\geq\frac{c}{[K(\alpha):K]}?$
+where $h_{S}$ is the standard height function (see Paragraph 4.1).
+In this paper, we show that Lehmer’s problem generalizes to a natural adelic curve defined on the field of rational functions with coefficients in $\mathbb{Q}$.
+Let us denote by $S$ the adelic curve on $\mathbb{Q}(T)$ defined in Paragraph 4.2.
+Our first main result is a generalization of Kronecker’s theorem to $S$.
+More precisely, we prove the following.
+Theorem 1.1 (A Kronecker’s theorem).
+$\{\alpha\in(\overline{\mathbb{Q}(T)})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
+Our second main result can be stated as follows.
+Theorem 1.2.
+Assuming Lehmer’s conjecture (see Conjecture 2.1 below), then for every $\alpha$ in an algebraic closure of $\mathbb{Q}(T)$ which is not a root of unity, we have
+$h_{S}(\alpha)\geq\frac{c}{\deg(\alpha)},$
+where $\deg(\alpha)=[\mathbb{Q}(T)(\alpha):\mathbb{Q}(T)]$ and $c$ is an absolute positive constant.
+(see Theorem 5.3).
+An interesting question is to explore the possibility of adapting Dobrowolski’s methods in our setting, which will be addressed in forthcoming papers.
+Let us note that generalizations of Lehmer’s problem to higher-dimensional varieties have been explored by Amoroso and David (cf.
+[1, 2, 3]).
+As a continuation of this work, we propose two conjectures, namely Conjectures 1.3 and 1.4 below, which we plan to investigate in a forthcoming paper.
+Let $n$ and $d$ be two positive integers.
+Conjecture 1.3.
+Let $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{G}_{m}^{n}\left(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})}\right)$, such that $\alpha_{1},\ldots,\alpha_{n}$ are multiplicatively independent.
+We have
+$h_{S}(\boldsymbol{\alpha})\geq\frac{c(n)}{\delta_{{\mathbb{Q}(T_{1},\ldots,T_{d})}}(\boldsymbol{\alpha})},$
+where $h_{S}$ is the height associated with the adelic curve $\mathbb{Q}(T_{1},\ldots,T_{d})$, and $\delta_{\mathbb{Q}(T_{1},\ldots,T_{d})}$ is, by definition, $\deg F_{0}$ where $F_{0}\in\mathbb{Q}(T_{1},\ldots,T_{d})[x_{1},\ldots,x_{n}]$ is a non-zero polynomial that vanishes at $\boldsymbol{\alpha}$ and has minimal degree.
+By Remark 5.2, it is clear that this constitutes a natural generalization of Conjecture 2.4 due to Amoroso and David.
+Observe that more refined versions have been developed, taking into account the possibility that $\alpha$ lies in a proper subgroup, see for instance, [16, 10, 11].
+Let $V$ be a hypersurface in $\mathbb{G}_{m}^{n}$ defined over $\mathbb{Q}(T_{1},\ldots,T_{d})$.
+Let us consider the normalized height $\hat{h}_{S}(V)$ of $V$, which can be defined using methods of [12].
+Conjecture 1.4.
+Assume that $V$ is defined over $\mathbb{Q}(T_{1},\ldots,T_{d})$ and that it is not an union of torsion subvarieties. Then,
+$\hat{h}_{S}(V)\geq c(n)\,.$
+where $c(n)$ is a positive constant which does not depend on $V$.
+Note that if $V$ is defined over $\mathbb{Q}$, then one can show that $\widehat{h}_{S}(V)=m(f)$, where $m(f)$ is the Mahler measure of an equation $f$ defining $V$ (cf.
+§2 for details).
+Conjecture 1.4 then reduces to Conjecture 2.3.
+Organization of the paper: In §2, we provide the definition of the usual Weil height function on $\mathbb{P}^{N}(\overline{\mathbb{Q}})$ and recall the classical Lehmer’s problem along with some results concerning this problem.
+In §3, we review the definition of the canonical height on Drinfel’d modules and briefly summarize some results concerning Lehmer’s problem in this context.
+In §4, we introduce the theory of adelic curves and discuss the extension of Lehmer’s problem to this setting.
+Finally, in §5, we prove our main results on Lehmer’s problem on the adelic curve $\mathbb{Q}(T)$.
+2. Lehmer’s problem on $\overline{\mathbb{Q}}$
+We denote by $\mathbb{Q}$ the field of rational numbers, $\mathbb{Z}$ the ring of rational integers, $K$ a number field, $\mathcal{O}_{K}$ its ring of integers, and $\overline{\mathbb{Q}}$ an algebraic closure of $\mathbb{Q}$.
+We denote by $\mathbb{R}$ the field of real numbers and $\mathbb{C}$ the field of complex numbers.
+We denote by $\mathbb{N}$ the set of natural numbers including 0.
+For each rational prime $p$, we denote by $|\cdot|_{p}$ the $p$-adic absolute value on $\mathbb{Q}$ such that $|p|_{p}=p^{-1}$; we also denote by $|\cdot|_{\infty}$ or simply $|\cdot|$ the standard absolute value.
+These form a complete set of absolute values on $\mathbb{Q}$: we identify the set $M_{\mathbb{Q}}$ of these absolute values with the set $\{\infty,p\mid p\text{ prime}\}$.
+More generally, we denote by $M_{K}$ the set of places of $K$ extending those of $M_{\mathbb{Q}}$.
+Let $x\in K$, we define the absolute value $|x|_{v}=|N_{K_{v}/\mathbb{Q}_{v}}(x)|_{v}^{1/[K_{\nu}:\mathbb{Q}_{v}]}$ for each $v\in M_{K}$, where $K_{v}:\mathbb{Q}_{v}]$ is the local degree at $v$.
+We have the product formula
+$\prod_{v\in M_{K}}|x|_{v}^{[K_{v}:\mathbb{Q}_{v}]}=1,$
+for all $x\in K\setminus\{0\}$ (cf. [22, Chapter V]).
+Let $K$ be a number field, $\mathcal{O}_{K}$ be the ring of integers, and $M_{K}$ be the set of places.
+Let $N$ be a positive integer.
+Let $\mathbb{P}^{n}$ be the projective space of dimension $N$.
+Let $P=[x_{0}:\cdots:x_{n}]\in\mathbb{P}^{N}(K)$.
+We define the absolute logarithmic height of the point $P$ as
+$h(P)=\sum_{v\in M_{K}}\frac{[K_{v}:\mathbb{Q}_{v}]}{[K:\mathbb{Q}]}\log\max(|x_{0}|_{v},\ldots,|x_{n}|_{v}),$
+which is independent of the choice of the projective coordinates by the product formula.
+The height $h(P)$ is independent of the choice of the base field $K$ (cf.
+[21, Chapter 3]).
+This defines the usual Weil height on $\mathbb{P}^{n}(\overline{\mathbb{Q}})$.
+Let $f\in\mathbb{C}[x_{1},\ldots,x_{n}]$. We define its Mahler measure by setting $m(0)=0$ and
+$m(f)=\frac{1}{(2\pi)^{n}}\int_{0}^{2\pi}\cdots\int_{0}^{2\pi}\log|f(e^{i\theta_{1}},\ldots,e^{i\theta_{n}})|\,d\theta_{1}\cdots d\theta_{n},$
+if $f\neq 0$. Let $f$ be a nonzero polynomial with complex coefficients in one variable such that $f(x)=a\prod_{i}(x-\alpha_{i})$. By Jensen’s formula, the Mahler measure of $f$ can be expressed as:
 (1) $m(f)=\log|a|+\sum_{i}\log\max(1,|\alpha_{i}|).$
-If $f$ has integer coefficients then (1) shows that $m(f)$ is nonnegative.
-A classical theorem of Kronecker asserts that $m(f)=0$ if and only if $|a|=1$ and $\alpha_{i}$ are zero or roots of the unity.
-Lehmer asked whether there exists a constant $c>1$ such that for every polynomial $f\in\mathbb{Z}[x]$ with $m(f)>0$, then $m(f)\geq c$.
-Lehmer’s question has not yet been settled.
-The best result in the direction of solving Lehmer’s problem is Dobrowolski’s result [10] which gives a lower bound depending on the degree of $f$.
-Let $h:\overline{\mathbb{Q}}\rightarrow\mathbb{R}$ be the absolute logarithmic height.
-Then, Lehmer’s problem can be reformulated as follows.
-Let $\alpha$ be an algebraic number of degree $d$ over $\mathbb{Q}$ and suppose that it is not a root of unity, then there exists an absolute constant $c>0$ such that
+If $f$ has integer coefficients, is irreducible, and has a positive degree $d$, then (1) shows that $m(f)$ is nonnegative, and
+$h(\alpha)=\frac{m(f)}{d},$
+where $\alpha$ is a root of $f$.
+A classical theorem of Kronecker asserts that when $f\in\mathbb{Z}[x]$, we have $m(f)=0$ if and only if $|a|=1$ and $\alpha_{i}$ are zero or roots of the unity.
+Boyd, Lawton, and Smyth independently showed (cf.
+[7, 23, 26]) that the Mahler measure of an irreducible polynomial
+$f\in\mathbb{Z}[x_{1},\ldots,x_{n}],$
+with $f\neq\pm x_{j}$, is equal to 0 if and only if the zero set of $f$ is a torsion variety.
+Let us point out that the case of hypersurfaces is particularly interesting.
+The notion of normalized height is linked to the Mahler measure of one of its equations.
+More precisely, let $V$ be a hypersurface of $\mathbb{G}_{m}^{n}$ defined over $\mathbb{Q}$ and let $f=0$ be one of its equations with integer coefficients of content 1.
+The normalized height of $V$ for the projective embedding that we have fixed, $\mathbb{G}_{m}^{n}\hookrightarrow\mathbb{P}^{n}$, is nothing other than the Mahler measure of the polynomial $f$:
+$\widehat{h}(V)=m(f).$
+(cf. [15, Proposition 2.1(vii)]).
+Lehmer asked whether there exists a constant $c>0$ such that for every polynomial $f\in\mathbb{Z}[x]$ with $m(f)>0$, then $m(f)\geq c$.
+Lehmer’s question can thus be translated into the following two equivalent conjectures:
+Conjecture 2.1.
+There exists a real number $c>0$ such that for every algebraic number $\alpha\in\overline{\mathbb{Q}}^{*}$ of degree $d$ over $\mathbb{Q}$ that is not a root of unity, we have
 $h(\alpha)\geq\frac{c}{d}.$
-Analogues of this conjecture have considered in other contexts, see [9, 8, 11].
-Let us point out that generalizations of Lehmer’s problem to higher dimensional varieties have been investigated by Amoroso and David, see for instance [1, 2, 3].
-Chen and Moriwaki [7] have introduced the theory of adelic curves.
-For example, a number field can be viewed as an example of adelic curves.
-More interestingly, they developed an arithmetic intersection theory on these curves.
-For instance, one can define a height function on the set of points of adelic curves.
-In this paper, we show that Lehmer’s problem generalizes to a natural adelic curve defined on the field of rational functions with coefficients in $\mathbb{Q}$.
-Let us denote by $S$ the adelic curve on $\mathbb{Q}(T_{1},\ldots,T_{d})$ defined in Section 2.2.
-In Theorem 4.1, we establish a generalization of Kronecker’s theorem to $S$.
-$\{\alpha\in(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
-where $h_{S}(\alpha)$ denotes is the height of the point.
-$[1:\alpha]$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})$, see Paragraph 2.1 or [12, Paragraph 3.2].
-$h_{S}$ is called naive height function in [12].
-Assuming Lehmer’s conjecture, then for every $\alpha$ in an algebraic closure of $\mathbb{Q}(T_{1},\ldots,T_{d})$ which is not a root of unity,
-$h_{S}(\alpha)\geq\frac{c}{\deg(\alpha)},$
-where $h_{S}$ is the standard height function (see (1)), $\deg(\alpha)=[\mathbb{Q}(T_{1},\ldots,T_{d})(\alpha):\mathbb{Q}(T_{1},\ldots,T_{d})]$ and $c$ is an absolute constant, see Theorem 4.3.
+Conjecture 2.2.
+There exists a real number $c>0$ such that for every irreducible polynomial $f\in\mathbb{Z}[x]$, $f\neq\pm x$, which is not a cyclotomic polynomial, we have
+$m(f)\geq c.$
+Within the framework of Conjectures 2.1 and 2.2, the best known result to date is Dobrowolski’s lower bound (see [18]): There exists a constant $c$ such that for every $\alpha\in\overline{\mathbb{Q}}^{*}$, an algebraic number of degree $d$, and is not a root of unity, we have
+(2) $h(\alpha)\geq\frac{c}{d}\left(\frac{\log(\log(3d))}{\log(3d)}\right)^{3}.$
+This inequality can obviously be reformulated in polynomial language:
+$m(f)\geq c\left(\frac{\log(\log(3d))}{\log(3d)}\right)^{3}$
+for any irreducible polynomial $f\in\mathbb{Z}[x]$, $f\neq\pm x$, of degree $d$, which is not a cyclotomic polynomial.
+Analogues of Conjectures 2.1 and 2.2 have been considered in the context of elliptic curves (cf. [13, 20]).
+Let us note that generalizations of Lehmer’s problem to higher-dimensional varieties have been explored by Amoroso and David (cf. [1, 2, 3]). For instance, they proposed the following conjectures.
+Conjecture 2.3.
+For every positive integer $n$, there exists a constant $c(n)>0$ such that for any irreducible polynomial $f\in\mathbb{Z}[x_{1},\ldots,x_{n}]$, $f\neq\pm x_{j}$, which is not a generalized cyclotomic polynomial, we have
+$m(f)\geq c(n).$
+Let $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{G}_{m}(\overline{\mathbb{Q}})^{n}$.
+Let $F_{0}\in\mathbb{Q}[x_{1},\ldots,x_{n}]$ be a non-zero polynomial that vanishes at $\boldsymbol{\alpha}$ and has minimal degree.
+Amoroso and David [2] defined the obstruction index of the point $\boldsymbol{\alpha}$ as the quantity $\deg F_{0}$, which they denote by $\delta(\boldsymbol{\alpha})$.
+Amoroso and David noted that
+$1\leq\delta(\boldsymbol{\alpha})\leq nD^{1/n},$
+where $D=[\mathbb{Q}(\alpha_{1},\ldots,\alpha_{n}):\mathbb{Q}]$.
+Conjecture 2.4.
+Let $n$ be a positive integer.
+There then exists a real number $c(n)>0$ such that for all $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{G}_{m}^{n}(\overline{\mathbb{Q}})$, such that $\alpha_{1},\ldots,\alpha_{n}$ are multiplicatively independent, we have
+$h(\boldsymbol{\alpha})\geq\frac{c(n)}{\delta(\boldsymbol{\alpha})}.$
+Amoroso and David [1] adapted the method of Dobrowolski in order to address Conjecture 2.4. Their work yielded the following result:
+$h(\boldsymbol{\alpha})\geqslant\frac{c(n)}{\delta(\boldsymbol{\alpha})}\log(3\delta(\boldsymbol{\alpha}))^{-\kappa(n)},$
+where $\kappa(n)=(n+1)(n+1)!^{n}-n$ (cf. [1, Théorème 1.5]).
+3. Lehmer’s problem for a Drinfel’d module
+Let $p$ be a prime number and $n$ a positive integer.
+We put $q:=p^{n}$, and we denote by $\mathbb{F}_{q}$ the finite field of cardinal $q$.
+We denote by $\mathbb{F}_{q}[T]$ the ring of polynomials in one variable with coefficients in the finite field $\mathbb{F}_{q}$, by $K=\mathbb{F}_{q}(T)$ its field of fractions, and by $K_{\infty}=\mathbb{F}_{q}((1/T))$ the completion of $K$ with respect to the $(1/T)$-adic valuation $v$.
+This valuation $v$ is extended to an algebraic closure $\overline{K}$ (respectively, $\overline{K}_{\infty}$) of $K$ (respectively, $K_{\infty}$).
+We also denote by $\iota$ the canonical homomorphism from $\mathbb{F}_{q}[T]$ to $\overline{K}$.
+We have the usual Weil height on $\mathbb{P}^{N}(\overline{K})$. If $L$ is a finite extension of $K$ of degree $m$, the height of a point $P=[x_{0}:\cdots:x_{N}]$ of $\mathbb{P}^{N}(L)$ is given by:
+$h(P)=\sum_{m}\frac{d(w)}{m}\max_{0\leq i\leq N}\left(-w(x_{i})\right),$
+where the sum is extended over all places of $L$, $d(w)$ denotes the residual degree at the place $w$, normalized by $w(L)=\mathbb{Z}\cup\{+\infty\}$.
+This height is independent of the field $L$ chosen as shown by the usual properties of valuations (cf.
+[21]).
+We define the height $h(\alpha)$ of an element $\alpha$ of $\overline{K}$ as the height $h([1:\alpha])$ of the point $1:\alpha]$.
+That is, we embed $\overline{K}$ into $\mathbb{P}^{1}$ and consider the Weil height relative to this embedding.
+Let $\overline{K}\{\tau\}$ be the ring of twisted polynomials over $\overline{K}$.
+For every $f:=\sum_{i=0}^{n}a_{i}\tau^{i}\in\overline{K}\{\tau\}$, we define its derivation $D$ at 0 by $D(f):=a_{0}$.
+An homomorphism of $\mathbb{F}_{q}$-algebras $\phi:\mathbb{F}_{q}[T]\to\overline{K}\{\tau\}$ is called a Drinfel’d $\mathbb{F}_{q}[T]$-module defined over $\overline{K}$ if $D\circ\phi=\iota$ and there exists $P\in\mathbb{F}_{q}[T]$ such that $\phi(P)\neq\iota(P)\tau^{0}$.
+According to [19, Chap.
+4, Lemma 4.5.1], there exists an integer $r>0$ such that $\deg(\phi(P))=r\deg(P)$, for all $P\in\mathbb{F}_{q}[T]$.
+The integer $r$ is called the rank of $\phi$.
+Denis [17] defined the global height of the Drinfel’d module $\phi$ at $\alpha\in\overline{K}$ as follows:
+$\hat{h}_{\phi}(\alpha)=\lim_{n\to\infty}\frac{h(\phi(P^{n})(\alpha))}{q^{nr}},$
+where $P$ is a polynomial in $\mathbb{F}_{q}[T]$ of positive degree $r$.
+This limit does not depend on the choice of $P$ (cf.
+[17, Théorème 1]).
+Furthermore, Denis [17, Théorème 2] established an analogue of Dobrowolski’s result [18] when $\phi$ is a Carlitz module.
+David and Pacheco [14] solved the so-called abelian version of the Lehmer’s problem in the context of Drinfel’d modules.
+It is worth noting that the classical version of this problem, concerning non-torsion points in $\mathbb{G}_{m}(\mathbb{Q}^{\text{ab}})$, was solved by Amoroso and Dvornicich [4].
+The analogue of this result for abelian varieties was given by Baker and Silverman, see [5, 6, 25].
 4. Adelic curves
@@
 Following [12], an adelic structure on $K$ consists of a measure space $(\Omega,\mathcal{A},\nu)$ equipped with a map $\phi:\omega\mapsto|\cdot|_{\omega}$ from $\Omega$ to $M_{K}$ satisfying the following properties:
-  1. (i)
+  1. (1)
     $\mathcal{A}$ is a $\sigma$-algebra on $\Omega$ and $\nu$ is a measure on $(\Omega,\mathcal{A})$;
-  2. (ii)
+  2. (2)
     for any $a\in K^{\ast}:=K\setminus\{0\}$, the function $\omega\mapsto\log|a|_{\omega}$ is $\mathcal{A}$-measurable, integrable with respect to the measure $\nu$.
@@
 holds for each $a\in K^{\ast}$.
+Number fields and function fields are fundamental examples of proper adelic curves (cf. [12, p. 170]). Note that the properness of these curves corresponds to the product formula.
 From now on, $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is a proper adelic curve and $K^{ac}$ an algebraic closure of $K$.
-Let $S\otimes_{K}K^{ac}=({K^{\mathrm{ac}}},(\Omega_{K^{\mathrm{ac}}},\mathcal{A}_{K^{\mathrm{ac}}},\nu_{K^{\mathrm{ac}}}),\phi_{K^{\mathrm{ac}}})$ be the algebraic extension of $S$ by ${K^{\mathrm{ac}}}$ (see [7, Paragraph 3.4.2] for the construction).
+Let $S\otimes_{K}K^{ac}=({K^{\mathrm{ac}}},(\Omega_{K^{\mathrm{ac}}},\mathcal{A}_{K^{\mathrm{ac}}},\nu_{K^{\mathrm{ac}}}),\phi_{K^{\mathrm{ac}}})$ be the algebraic extension of $S$ by $K^{\mathrm{ac}}$} (see [12, Paragraph 3.4.2] for the construction).
 4.1. Height of points in $\mathbb{P}^{n}(K^{\mathrm{ac}})$
-For $[a_{0}:\ldots:a_{n}]\in\mathbb{P}^{n}({K^{\mathrm{ac}}})$, the height of $[a_{0}:\ldots:a_{n}]$ with respect to the adelic curve $S$ is given as follows.
+For $a_{0}:\ldots:a_{n}]\in\mathbb{P}^{n}({K^{\mathrm{ac}}})$, the height of $a_{0}:\ldots:a_{n}]$ with respect to the adelic curve $S$ is given as follows.
 $h_{S}([a_{0}:\ldots:a_{n}])=\int_{\Omega^{ac}}\log\max(|a_{0}|_{\chi},\ldots,|a_{n}|_{\chi})\nu_{K^{\mathrm{ac}}}(d\chi),$
-see [7, Paragraph 3.5] for more details about the construction.
+see [12, Paragraph 3.5] for more details about the construction.
 For every $\alpha\in K^{\textrm{ac}}$, we set
@@
 Let $K=\mathbb{Q}(T_{1},\ldots,T_{d})$ be the field of rational functions of $d$ variables $T_{1},\ldots,T_{d}$ and with coefficients in $\mathbb{Q}$.
-$\mathbb{Q}(T_{1},\ldots,T_{d})$ can be endowed with an interesting adelic structure, see [7, Paragraph 3.2.6], or [12, Section 3] and [6, Example 2.5] for a more explicit construction.
-Let us review in details the case when $d=1$.
-Let $\Omega=\Omega_{h}\coprod\mathcal{P}\coprod[0,1]_{\ast}$, where $\Omega_{h}$ is the set of all closed points of $\mathbb{P}_{\mathbb{Q}}^{1}\setminus\{\infty\}$, and $\mathcal{P}$ denotes the set all prime numbers, and $[0,1]_{\ast}$ the set of $t\in[0,1]$ such that $e^{2\pi it}$ is transcendental.
+$\mathbb{Q}(T_{1},\ldots,T_{d})$ can be endowed with an interesting adelic structure, see [12, Paragraph 3.2.6], or [24, Section 3] and [9, Example 2.5] for a more explicit construction.
+Let us review in detail the case when $d=1$.
+Let $\Omega=\Omega_{h}\coprod\mathcal{P}\coprod[0,1]_{\ast}$, where $\Omega_{h}$ is the set of all closed points of $\mathbb{P}_{\mathbb{Q}}^{1}\setminus\{\infty\}$, and $\mathcal{P}$ denotes the set all prime numbers, and $0,1]_{\ast}$ the set of $t\in[0,1]$ such that $e^{2\pi it}$ is transcendental.
 Let $\phi:\Omega\rightarrow M_{K}$ be the map sending $\omega\in\Omega$ to $|\cdot|_{\omega}$, defined as follows.
 Let $x\in\Omega_{h}$.
-$x$ defines a discrete valuation on $K$ which is denoted by $\mathrm{ord}_{x}(\cdot)$.
+Then, $x$ defines a discrete valuation on $K$ which is denoted by $\mathrm{ord}_{x}(\cdot)$.
 Let $F_{x}$ be an irreducible polynomial in $\mathbb{Z}[T]$ with $F_{x}(x)=0$.
@@
 $|\varphi|_{t}:=|\varphi(e^{2\pi it})|,\quad\forall\varphi\in\mathbb{Q}(T),$
-where ||$\cdot$|| ⋅ | denotes the usual absolute value of $\mathbb{C}$.
-$\Omega_{h}$ and $\mathcal{P}$ are equipped with the discrete $\sigma$-algebras, and $[0,1]_{\ast}$ with the restriction of the Borel $\sigma$-algebra on $[0,1]$.
-Let $\mathcal{A}$ be the $\sigma$-algebra on $\Omega$ generated by the above $\sigma$-algebras on $\Omega_{h},\mathcal{P}$ and $[0,1]_{\ast}$ respectively.
-Let $\nu$ be the measure on $\Omega$ such that $\nu(\{x\})=1$ for $x\in\Omega_{h}$, that $\nu(\{p\})=1$ for any prime number $p$ and that the restriction of $\nu$ on $[0,1]_{\ast}$ coincides with the Lebesgue measure.
-This data defines an adelic structure on $\mathbb{Q}(T)$. Moreover, $(\mathbb{Q}(T),(\Omega,\mathcal{A},\nu),\phi)$ is proper, see [7, Paragraph 3.2.5].
-3. On Lehmer’s problem on adelic curves
-Question 3.1.
+where $|\cdot|$ denotes the usual absolute value of $\mathbb{C}$.
+$\Omega_{h}$ and $\mathcal{P}$ are equipped with the discrete $\sigma$-algebras, and $0,1]_{\ast}$ with the restriction of the Borel $\sigma$-algebra on $[0,1]$.
+Let $\mathcal{A}$ be the $\sigma$-algebra on $\Omega$ generated by the above $\sigma$-algebras on $\Omega_{h},\mathcal{P}$ and $0,1]_{\ast}$ respectively.
+Let $\nu$ be the measure on $\Omega$ such that $\nu(\{x\})=1$ for $x\in\Omega_{h}$, that $\nu(\{p\})=1$ for any prime number $p$ and that the restriction of $\nu$ on $0,1]_{\ast}$ coincides with the Lebesgue measure.
+This data defines an adelic structure on $\mathbb{Q}(T)$. Moreover, $(\mathbb{Q}(T),(\Omega,\mathcal{A},\nu),\phi)$ is proper (cf. [12, Paragraph 3.2.5]).
+4.3. On Lehmer’s problem on adelic curves
+Motivated by the previous work on Lehmer’s problem, it is natural to consider the following question:
+Question 4.2.
 Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve and $K^{ac}$ an algebraic closure of $K$.
-Is there a constant $c>0$ such for every $\alpha\in K^{\mathrm{ac}}$ with $h_{S}(\alpha)>0$, then
-$h_{S}(\alpha)\geq\frac{c}{[K(\alpha):K]}.$
+Is there a constant $c:=c_{S}>0$ such for every $\alpha\in K^{\mathrm{ac}}$ with $h_{S}(\alpha)>0$, we have
+$h_{S}(\alpha)\geq\frac{c}{[K(\alpha):K]}?$
 When $K$ is $\mathbb{Q}$ and $S$ is the classical adelic curve associated with $\mathbb{Q}$, then this is the classical Lehmer’s problem.
-Next, we show that the classical Lehmer’s problem generalizes naturally to the adelic curve $\mathbb{Q}(T)$.
-4. Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$
-Let $S$ be the adelic curve on $\mathbb{Q}(T)$ recalled in Paragraph 2.2. Let
-$[\varphi:\psi]\in\mathbb{P}^{1}\left(\overline{\mathbb{Q}(T)}\right).$
+5. Lehmer’s problem on $\overline{\mathbb{Q}(T)}$
+In this section, we show that the classical Lehmer’s problem generalizes naturally to the adelic curve $\mathbb{Q}(T)$.
+Let $S$ be the adelic curve on $\mathbb{Q}(T)$ recalled in Paragraph 4.2. Let
+$\varphi:\psi]\in\mathbb{P}^{1}\left(\overline{\mathbb{Q}(T)}\right).$
 We may assume that $\varphi$ and $\psi$ are integral over $\mathbb{Z}[T]$.
@@
 $\frac{1}{[K^{\prime}:\mathbb{Q}(T)]}\max\Bigl(\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr|_{\omega}dt,\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr|_{\omega}dt\Bigr)\leq h_{S}([\varphi:\psi]).$
-Using [7, Lemma 3.5.4], we obtain
+Using [12, Lemma 3.5.4], we obtain
 $\begin{split}\tfrac{1}{[K^{\prime}:\mathbb{Q}(T)]}\max\Bigl(\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl|\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr|_{\omega}dt,\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}&\log\Bigl|\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr|_{\omega}dt\Bigr)\\ &\leq[K^{\prime}:\mathbb{Q}(T)]h_{S}([\varphi:\psi]).\end{split}$
@@
 (4) $\begin{split}\max\Bigl(\int_{t\in[0,1]}\log\Bigl|\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr|_{t}dt,\int_{t\in[0,1]}&\log\Bigl|\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr|_{t}dt\Bigr)\\ &\leq[K^{\prime}:\mathbb{Q}(T)]h_{S}([\varphi:\psi]).\end{split}$
-Theorem 4.1.
-(A Kronecker’s theorem)
-$\{\alpha\in(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
+Theorem 5.1 (A Kronecker’s theorem).
+$\{\alpha\in(\overline{\mathbb{Q}(T)})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
 Proof.
-For simplicity, we prove the theorem when $d=1$. Let $\varphi$ and $\psi$ be two coprime polynomials in $\mathbb{Z}[T]$. We have
+Let $\varphi$ and $\psi$ be two coprime polynomials in $\mathbb{Z}[T]$. We have
 $h_{S}([\varphi:\psi])=\int_{0}^{1}\log\max(|\varphi(e^{2\pi it})|,|\psi(e^{2\pi it})|)dt.$
@@
 Let $\varphi\in\overline{\mathbb{Q}(T)}$. Let $d:=[\mathbb{Q}(T)(\varphi):\mathbb{Q}(T)]$. We can assume
-$[1:\varphi]=[\varphi_{1}:\varphi_{2}],$
+$1:\varphi]=[\varphi_{1}:\varphi_{2}],$
 with $\varphi_{1}\in\mathbb{Z}[T]$ and $\varphi_{2}\in\overline{\mathbb{Z}[T]}$ (integral closure of $\mathbb{Z}[T]$).
@@
 Remark 5.2.
-If $\alpha\in\overline{\mathbb{Q}}$, then
+Let $\alpha\in\overline{\mathbb{Q}}$. It is clear that
 $h_{S}(\alpha)=h(\alpha),$
-where $h$ is the standard Weil’s height of $\alpha$.
+where $h$ is the standard Weil height of $\alpha$.
 We conclude that
 Theorem 5.3.
-Lehmer’s problem is equivalent to Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$.
+Lehmer’s problem is equivalent to Lehmer’s problem on $\mathbb{Q}(T)$.
+Acknowledgement: I would like to thank the referee for their insightful comments and suggestions.